### v1

# CancerGUIDE: Cancer Guideline Understanding via Internal Disagreement Estimation

Abstract
The National Comprehensive Cancer Network (NCCN) provides evidence-based guidelines for cancer treatment. Translating complex patient presentations into guideline-compliant treatment recommendations is time-intensive, requires specialized expertise, and is prone to error. Advances in large language model (LLM) capabilities promise to reduce the time required to generate treatment recommendations and improve accuracy. We present an LLM agent-based approach to automatically generate guideline-concordant treatment trajectories for patients with non-small cell lung cancer (NSCLC). Our contributions are threefold. First, we construct a novel longitudinal dataset of 121 cases of NSCLC patients that includes clinical encounters, diagnostic results, and medical histories, each expertly annotated with the corresponding NCCN guideline trajectories by board-certified oncologists. Second, we demonstrate that existing LLMs possess domain-specific knowledge that enables high-quality proxy benchmark generation for both model development and evaluation, achieving strong correlation (Spearman coefficient $r=0.88$, RMSE = $0.08$) with expert-annotated benchmarks. Third, we develop a hybrid approach combining expensive human annotations with model consistency information to create both the agent framework that predicts the relevant guidelines for a patient, as well as a meta-classifier that verifies prediction accuracy with calibrated confidence scores for treatment recommendations (AUROC=0.800), a critical capability for communicating the accuracy of outputs, custom-tailoring tradeoffs in performance, and supporting regulatory compliance. This work establishes a framework for clinically viable LLM-based guideline adherence systems that balance accuracy, interpretability, and regulatory requirements while reducing annotation costs, providing a scalable pathway toward automated clinical decision support.
keywords:
Treatment recommendation, Cancer, Guidelines, Evaluation, Scaling
Data and Code Availability
We obtain clinical annotations from a third-party contractor. Patient data is de-identified and sourced from multiple institutions within the USA, but it is not licensed for public release. Code is made available here: CancerGUIDE repository.
Institutional Review Board (IRB)
Our research does not contain data from identified human subjects and as such does not require an IRB.
1 Introduction
Cancer treatment decisions require oncologists to synthesize complex patient histories with evolving clinical guidelines to recommend the appropriate next treatments. The National Comprehensive Cancer Network (NCCN) guidelines provide evidence- and consensus-based recommendations for cancer diagnosis, treatment, and management (National Comprehensive Cancer Network, 2025), with adherence promoting higher quality and consistent care between providers (Pluchino and D’Amico, 2020; Winn and McClure, 2003; Benson and Brown, 2008). However, guideline navigation presents significant challenges: the guidelines are extensive, frequently updated as new research emerges, and require time-intensive review of complex patient documentation (Benson and Brown, 2008). These factors contribute to variability in guideline adherence and treatment recommendations, particularly in resource-constrained settings where specialist expertise is limited (Koh et al., 2020; Kumar et al., 2023).
Large language models (LLMs) offer promising potential to address these challenges by automatically processing clinical notes and recommending guideline-concordant treatment plans (Fast et al., 2024; Wang et al., 2024; Hamed et al., 2023; Schulte, 2023). However, deploying LLMs for clinical decision support requires rigorous evaluation to ensure accuracy and safety. The complexity of clinical reasoning, combined with the high stakes of treatment decisions, demands robust validation methods that can assess model performance at scale while maintaining clinical safety standards. In addition, current FDA guidelines for evaluation of AI systems recommend ROC curve measurement as a part of comprehensive clinical performance assessment endpoints (U.S. Food and Drug Administration, 2022), which is challenging to produce from generative outputs that are not associated with semantically aligned confidence scores.
Despite growing interest in clinical LLMs, rigorous evaluation remains a fundamental bottleneck due to the scarcity of expert-annotated datasets. High-quality ground truth labels for complex clinical reasoning tasks require substantial investment in specialist time and expertise, limiting the scale at which models can be validated (Zajac et al., 2023; Liang et al., 2017). Common evaluation approaches face significant limitations: synthetic data generation often fails to capture clinical complexity and is vulnerable to distributional shift (Santangelo et al., 2025; Yan et al., 2022; Chen et al., 2024; Cui et al., 2025), while using actual patient treatments as ground truth is problematic, since real-world decisions frequently incorporate factors exogenous to guideline recommendations, such as patient preferences, drug availability, institutional protocols, and physician experience (Arts et al., 2016; Quaglini et al., 2004).
This evaluation challenge is particularly acute for guideline adherence tasks, where the gold standard requires expert oncologists to determine whether complex, multi-step clinical reasoning aligns with evidence-based recommendations. The resulting annotation bottleneck creates a critical gap: while LLMs show promise for clinical decision support, practitioners lack scalable methods to assess model reliability before deployment in high-stakes healthcare settings.
Figure 1: CancerGUIDE framework: (1) Clinicians derive patient pathways ($label_{doc}$) from NCCN guidelines and real notes, which are then compared with LLM-derived pathways ($label_{LLM}$) to calculate $Score_{r}$. (2) Proxy benchmarks use NCCN guidelines and both synthetic and real clinical notes to evaluate LLMs by comparing their outputs ($label_{LLM}$) against weak labels ($label_{proxy}$) to calculate $Score_{p}$ and measure similarity of performance on expert annotations. (3) A consistency framework uses model consistency features (Feature Sets) and $Score_{p}$ in a meta-classifier to decide whether to accept or reject treatment predictions.
We address this evaluation bottleneck through two complementary approaches that enable scalable assessment of LLM performance on guideline adherence without extensive expert annotation. First, we evaluate models across six different proxy-benchmark generation methods: two synthetically generated datasets and four based on real clinical notes with consistency-derived labels. This analysis enables model selection and preliminary capability assessment without ground-truth labels. Second, we demonstrate that self- and cross-model consistency (the degree to which a model agrees with itself and other models) serve as reliable predictors of accuracy on expert-annotated cases.
To validate these approaches, we construct the first benchmark for NCCN guideline adherence on non-small cell lung cancer (NSCLC) by eliciting 13 oncologists to annotate 121 complete patient pathways, representing 130+ hours of specialist expertise. This expert-validated dataset enables us to systematically evaluate six different proxy benchmarking methods and quantify the relationship between model consistency and clinical accuracy across multiple frontier LLMs. We then develop a meta-classifier framework that combines these weak supervision signals to classify individual prediction correctness, achieving robust performance while requiring minimal expert validation (Figure 1).
Our evaluation highlights the role of synthetic data and consistency-based data in proxy benchmarking, with these approaches achieving high Spearman correlation coefficients (r=0.88) and low RMSE values (RMSE=0.08). We demonstrate that model consistency can also serve as a reliable accuracy predictor, enabling our meta-learning framework to achieve an average of 0.800 AUROC in classifying individual prediction accuracy across all models, while unsupervised clustering using consistency signals alone achieves 0.666 F1 at separating correct from incorrect predictions.
Our primary contributions are as follows:
  1. 1.
    NCCN guideline-adherent dataset and task formalization: The first rigorous ML problem formulation for clinical guideline adherence, with an expert-annotated dataset of 121 patient pathways and benchmarking across eight frontier LLMs.
  2. 2.
    Proxy benchmark to validate performance in zero-label settings: A systematic evaluation that identifies which proxy methods best predict clinical performance without expert labels.
  3. 3.
    Consistency framework for reliable treatment prediction: A hybrid agent and meta-classifier framework that generates relevant guideline paths and produces confidence values correlated with accuracy, enabling ROC curve calculation to clearly convey performance to clinicians and ensure compliance with regulatory standards (U.S. Food and Drug Administration, 2022).
2 Related Work
2.1 Guideline Adherent Treatment Recommendations
Guideline-adherent treatments have consistently been associated with improved clinical outcomes, such as overall survival (Dholakia et al., 2020; Lindqvist et al., 2022; Wöckel et al., 2010). Clinical decision-support systems have the ability to translate clinical guidelines into point-of-care recommendations. Several studies have demonstrated the use of AI-driven systems in predicting guideline-adherent treatment recommendations in oncology (Nafees et al., 2023; Suwanvecho et al., 2021). Emerging work has also explored the utility of LLMs; for example, preliminary investigations suggest that ChatGPT can effectively summarize guideline content (Hamed et al., 2023) and exhibits partial concordance in identifying guideline-adherent treatments (Schulte, 2023). Furthermore, recent benchmarking efforts have systematically evaluated the alignment of LLMs with established medical guidelines, underscoring both their potential utility and the current limitations imposed by the relatively small number of verified clinical cases available (Fast et al., 2024).
2.2 Synthetic Generation of Clinical Data
Synthetic data generation is a promising strategy to address the challenges of privacy and data scarcity in healthcare. Past works have focused on generating high-fidelity electronic health records (EHRs) using generative models, such as generative adversarial networks, variational autoencoders and autoregressive models (Choi et al., 2017; Lu et al., 2022; Li et al., 2023; Yoon et al., 2023; Jarrett et al., 2021). More recent frameworks have used LLMs to generate synthetic data and have shown improvements in privacy guarantees and scalability (Altalla’ et al., 2025; Barr et al., 2025; Taloni et al., 2025; Litake et al., 2024). In parallel, standardized evaluation metrics and scorecards have been proposed to assess the fidelity, privacy, and clinical utility of synthetic health data, providing a foundation for more rigorous benchmarking and deployment (Santangelo et al., 2025; Yan et al., 2022; Chen et al., 2024; Bedi et al., 2025). In addition, synthetic cohorts have been developed to model guideline-adherent treatment pathways, such as synthetic stroke registries for adherence benchmarking (Wang et al., 2025) and synthetic EHR modules that embed clinical practice guidelines and protocols (Walonoski et al., 2017). These efforts demonstrate the potential of synthetic data for adherence-focused tasks. However, significant challenges remain in generating high-fidelity, guideline-adherent cases while maintaining performance on downstream evaluation tasks.
2.3 Weak Supervision and Consistency
Weak supervision and consistency-driven learning have emerged as key strategies for training models when labeled data is limited or noisy. Weak supervision techniques, such as programmatic labeling and distant supervision, integrate heterogeneous, partially labeled sources to produce probabilistic labels for downstream predictive tasks (Mintz et al., 2009; Ratner et al., 2016, 2017; Hsu and Roberts, 2025; Saad-Falcon et al., 2025). Consistency-based methods encourage models to produce stable predictions under input perturbations, data augmentations, or repeated evaluations, helping to regularize training and improve generalization (Laine and Aila, 2016; Miyato et al., 2019; Xie et al., 2020). Pseudo-labeling, a related strategy, generates labels for unlabeled or partially labeled data by treating confident model predictions as additional supervised training data, enhancing model performance by expanding the training set with its own high-confidence predictions (Lee, 2013; Zou et al., 2021; Xie et al., 2020). Within healthcare, pseudo-labeling has shown to improve the reliability of EHR phenotyping and imaging tasks under label scarcity (Dong et al., 2023; Zhang et al., 2022; Nogues et al., 2022). Together, these techniques provide a framework for leveraging limited labeled data while maintaining prediction stability and adherence to domain-specific constraints, which is particularly relevant for guideline-adherent treatment modeling.
2.4 Non-Verifiable Task Evaluation
Evaluating LLMs on open-ended or non-verifiable tasks, such as treatment recommendation prediction, is challenging due to the scarcity of ground truth labels. Traditional metrics like accuracy or BLEU are insufficient in zero-label settings, prompting the use of human-aligned and/or proxy evaluation frameworks (Lyu et al., 2024; Yamauchi et al., 2025; Tan et al., 2024). Further, studies have shown that consistency-derived benchmarks provide improved predictive fidelity relative to ground-truth outcomes (Lee et al., 2024; Li et al., 2025). Proxy frameworks and consistency-derived benchmarks offer partial solutions, but their alignment with expert judgment remains imperfect, motivating the continued development of hybrid evaluation strategies.
3 Methods
3.1 Guideline-Compliance Task Formalization
We formalize guideline-compliant treatment prediction as a structured prediction problem. Let $x\in\mathcal{X}$ denote patient notes and $y\in\mathcal{Y}$ the corresponding guideline-compliant pathway, where $\mathcal{Y}$ is the space of decision graphs with nodes as clinical decisions and terminal nodes as treatments. An LLM $f:\mathcal{X}\to\mathcal{Y}$ produces predictions $\hat{y}=f(x)$.
To ground this formalization in clinical practice, we leveraged o3’s vision capabilities (OpenAI, 2025c) to extract the NCCN NSCLC guideline decision tree (National Comprehensive Cancer Network, 2025) and curated an expert-annotated dataset. Oncologists traced patient notes through the decision tree, recording node sequences and assessing guideline adherence for each case. Further details on annotation procedures, participant recruitment, and quality control measures are provided in Appendix A.
Since gold-standard pairs $(x,y)$ are difficult to obtain, direct supervision is not scalable as $X$ grows. To address this, we first introduce proxy benchmarking methods that act as substitutes for direct supervision. We define synthetic inputs $\hat{x}$ and corresponding predicted pathways $\hat{y}$, forming two primary classes of proxy evaluations:
  1. 1.
[…]
  2. 2.
[…]
These proxy-based evaluations provide measurable signals of model performance, enabling ranking of models or detection of likely errors in zero-label settings. We learn a surrogate evaluator $g$ that predicts whether $\hat{y}$ is guideline-compliant. Concretely, we define a feature mapping
$\phi:(\mathcal{X},\mathcal{Y},f)\to\mathbb{R}^{d}$
that extracts signals such as (i) model self-consistency across rollouts, (ii) agreement across models, and (iii) alignment with proxy benchmarks. The evaluator is then trained as a binary classifier
$g(\phi(x,\hat{y},f))\approx\mathbf{1}\{\hat{y}=y\}.$
Our objective is to minimize the expected classification loss
$\min_{g}\;\mathbb{E}_{(x,y)\sim\mathcal{D}}\big{[}\mathcal{L}(g(\phi(x,\hat{y},f)),\mathbf{1}\{\hat{y}=y\})\big{]},$
where $\mathcal{D}$ denotes the distribution over patients and $\mathcal{L}$ is standard loss. This formulation enables evaluation of $f$ in settings with limited or no access to human-labeled $(x,y)$ pairs, by leveraging mode agreement, self-consistency, and benchmark proxy-derived features as signals for meta-classification.
3.1.1 Evaluation Preliminaries
To evaluate model performance, we employ two complementary metrics:
  1. 1.
    Path Overlap: Measures the proportion of nodes in predicted paths that are repeated, relative to the total nodes in all compared paths. This captures consistency of decision sequences with the full derivation in Appendix B
  2. 2.
    Treatment Match: A binary score indicating whether the final predicted treatment matches the ground truth when available. In settings without ground truth, it is computed as the proportion of repeated final treatments across multiple predictions, with full derivation in Appendix C.
3.2 Zero-Label Benchmark Generation
To approximate $(x,y)$ pairs for evaluation, we introduce two complementary approaches (synthetic supervision and consistency-based pseudo-labeling), with a total of six proxy benchmarking methods for zero-label performance estimation contributed.
3.2.1 Synthetic Supervision
We generate high-fidelity synthetic patient notes $\hat{x}$ paired with generated guideline paths $\hat{y}$ to simulate real clinical cases. The goal is to maximize fidelity to realistic patient notes while ensuring the target guideline path is accurate. Our multi-step pipeline separates the generation of $\hat{x}$ and the selection of $\hat{y}$ to filter incorrect labels from the benchmark, while still maintaining meaningfully realistic patient cases.
Generating $\hat{x}$
Two complementary strategies are used: Structured Generation fills empty structured fields conditioned on the generated path and full clinical guidelines, performs consistency checks by reconstructing implied paths (discarding mismatched cases), then generates unstructured notes from structured data, target path, guidelines, and real clinical note examples. Unstructured Generation bypasses structured fields, generating synthetic notes directly from target paths, guidelines, and clinical note examples.
Selecting $\hat{y}$ via LLM Preference
Once $\hat{x}$ is generated, we obtain $\hat{y}$ by having the LLM generate the path from $\hat{x}$. If the prediction matches the target path, we accept the pair directly. Otherwise, the LLM chooses between predicted and target paths in position-agnostic format, and we retain only cases where the target path is selected.
[…]
3.2.2 Consistency-Based Pseudo-Labeling
We propose a consistency-based pseudo-labeling strategy to construct proxy benchmarks that approximate model performance on treatment prediction and guideline path generation. Consistency-based pseudo-labels are derived from two sources: (i) Self-consistency, which leverages agreement within repeated predictions of a single model, and (ii) Cross-Model Consistency, which relies on agreement across different models. For each source, we define two benchmarks based on whether consistency is measured with respect to the path overlap metric or the treatment match metric, yielding four benchmarks in total.
Self-Consistency
For real clinical notes, pseudo-labels are assigned using model self-consistency:
  1. 1.
    Sample $k$ independent predictions $f(x)=\{\hat{y}_{1},\dots,\hat{y}_{k}\}$ from the same model.
  2. 2.
    Measure agreement along two axes: path overlap (structural alignment) and treatment match (final treatment concordance).
  3. 3.
    Retain notes with agreement above a threshold $\delta=0.9$; assign inconsistent cases score of 0.
Cross-Model Consistency
We further assign pseudo-labels by comparing predictions across models $M$. For a note $x$, if two or more models converge on the same pseudo-label after $k$ independent samples, we define this as the aggregated label. Inconsistent cases are simply excluded rather than penalized.
3.3 Final Treatment Accuracy Prediction
Supervised Accuracy Classification
We use consistency signals to predict accuracy and identify errors without human labels. We train a meta-classifier using features derived from self-consistency metrics ($k$-rollout path overlap and treatment match), cross-model consistency (fraction of models with the same generated path), and proxy benchmark scores (synthetic and consistency-derived). This framework achieves an average of 0.800 AUROC in predicting whether the output of a model is correct, providing practitioners with a means of assessing confidence in model recommendations.
Unsupervised Performance for Accuracy Classification and Error Identification
To demonstrate label-free evaluation capabilities, we apply unsupervised methods to both accuracy classification and error identification. Clustering self- and cross-model consistency features naturally separates high-confidence correct from low-confidence incorrect predictions, achieving 0.666 F1 score without any human annotations. For error identification, we tabulate inconsistencies across $k$ rollouts for each model $m$ on patient $p$ and find that 40.42% of all model errors on human-labeled data can be detected without any human labels (detailed analysis in Appendix D). This demonstrates that consistency signals alone provide sufficient structure to both differentiate between accurate and inaccurate predictions and identify specific failure modes, enabling model developers to detect and address potential issues prior to deployment for iterative refinement of clinical decision support systems.
4 Results
4.1 Expert-Annotated Dataset
To evaluate LLM performance on guideline-concordant treatment prediction, we constructed a high-quality expert-annotated dataset.
Annotation Details
A total of 13 oncologists annotated 121 patient notes. To assess inter-annotator reliability, 11 examples were dually annotated, showing an average of 0.636 treatment match and 0.692 path overlap score. Notes were on average 54755.3 characters long and contained 82 distinct paths through the NCCN decision tree and 48 distinct final treatment recommendations. We use this annotation as ground truth for both evaluating model performance and validating proxy benchmarking approaches. By establishing a high-quality reference standard, we can interpret LLM performance, detect systematic errors, and calibrate models in high-stakes medical settings without relying exclusively on expensive ongoing human annotation.
Model Performance on Human-Annotated Benchmark
Table 1 illustrates the results of eight frontier models on the expert-annotated dataset. We report performance on GPT-5 (OpenAI, 2025b), GPT-4.1 (OpenAI, 2025a), o3 (OpenAI, 2025c), o4-mini (OpenAI, 2025d), DeepSeek-R1 (DeepSeek-AI, 2025), and LLaMA-3.3-70B-Instruct (Meta, 2024). We evaluate GPT-5 with varying reasoning efforts (minimal, medium, and high) to directly assess the impact of reasoning on performance. All models are evaluated with default temperature 1.0. We measured two clinically relevant metrics: (i) path overlap, measuring structural agreement with the annotated guideline path, and (ii) treatment match, indicating whether the recommended treatment node matches expert annotation.
Table 1: Model performance on Expert-Annotated Data
Model Path Overlap Treatment Match
GPT-5-High 0.455 $\pm$ 0.035 0.339 $\pm$ 0.043
GPT-5-Medium 0.483 $\pm$ 0.036 0.364 $\pm$ 0.044
GPT-5-Minimal 0.441 $\pm$ 0.032 0.322 $\pm$ 0.043
GPT-4.1 0.388 $\pm$ 0.035 0.298 $\pm$ 0.042
o3 0.477 $\pm$ 0.038 0.364 $\pm$ 0.044
o4-mini 0.433 $\pm$ 0.037 0.339 $\pm$ 0.043
Deepseek-R1 0.419 $\pm$ 0.038 0.355 $\pm$ 0.044
LLaMA-3.3-70B-Instr. 0.174 $\pm$ 0.022 0.112 $\pm$ 0.029
4.2 Proxy Benchmarks Enable Model Evaluation in Zero-Label Settings
We evaluated six proxy benchmarks for assessing LLM performance without human labels (Figure 2). Among synthetic approaches, introducing structure into the generation pipeline substantially increased error—RMSE rose from 0.11 (Synthetic Unstructured) to 0.43 (Synthetic Structured)—while providing only a negligible correlation gain (0.86 → 0.90). For consistency-based methods, thresholding by treatment match outperformed path overlap, likely because treatment match captures a more generalizable prediction goal, whereas path overlap harshly penalizes minor deviations with little impact on downstream treatment prediction. Aggregating outputs across models in cross-model consistency did not improve correlations and instead increased RMSE. Overall, synthetic benchmarks appear promising due to their scalability and robustness to variance between model consistency levels, but in domains where accuracy aligns with model consistency, consistency-based benchmarks can also provide a strong signal of performance. Examining model self-consistency in more detail reveals that outputs with higher internal agreement tend to be more accurate. We visualize this trend of iteration consistency in Figure 3 and report per-model Pearson correlation coefficients for both path overlap and treatment match in the appendix (Table 3), confirming that consistency is a strong indicator of instance-level accuracy.
Figure 2: Self-consistency pseudo-labels provide a robust proxy for benchmarking. Six approaches are evaluated: synthetic data (structured and unstructured), self-consistency pseudo-labeling (with varying acceptance criteria), and cross-model consistency pseudo-labeling. Correlation is measured using Spearman coefficients as well as root mean-squared error with color intensity indicating magnitude.
Figure 3: Model accuracy increases with self-consistency across prediction runs. Higher consistency (fraction of runs producing identical paths) correlates with improved performance for treatment matching.
4.3 Predicting Treatment Accuracy Using Model Self-Consistency
We develop a meta-classifier to predict when treatment recommendations are correct, using features from model consistency patterns and benchmark performance. Our approach achieves 0.800 AUROC by leveraging inference-time signals to predict accuracy without requiring ground truth labels at test time. We evaluate five feature sets to identify which signals contribute most to performance (Table 2). We stratify train and test sets by models and patient IDs with a 70/30 split. Performance generalizes across models and patients, enabling downstream classification with minimal human supervision.
Table 2: Feature sets for predicting treatment recommendation accuracy
Feature Set Self- Consistency Synth. Bench Cons. Bench Cross- Consistency
Base
Internal
Agg.
Base+Agg.
All
  • Agg. = Aggregated.
  • Self-consistency and Cross-Consistency features are computed per sample.
  • Synth. Bench = Synthetic Benchmarks; Cons. Bench = Consistency Benchmarks.
We find in Figure LABEL:fig:roc_aggregated that Base_aggregated, Aggregated_only, and All outperform other feature sets, indicating that cross-model consistency provides the strongest signal for accuracy classification. We see minimal gain from proxy benchmark performance, indicating consistency-based approaches are better suited for real-time analysis of model outputs than synthetic data. Even in settings without cross-model information (Base and Internal), we see high AUROC scores, indicating model self-consistency can be a strong signal for classification in isolation. Figure LABEL:fig:roc_base_agg_models shows intra-model variability of the classifier trained using Base_aggregated features, with results using Internal features reported in Appendix G. The strong performance of LLaMA-3.3-70B-Instruct highlights a key issue with including lower-quality models: its performance on the NCCN task is lower than other models, arbitrarily inflating AUROC as the confidence prediction task of the meta-classifier becomes easier. Besides LLaMA-3.3-70B-Instruct, GPT-4.1 has the highest AUROC and Deepseek-R1 the lowest, indicating GPT-4.1’s consistency better correlates with accuracy than Deepseek-R1.
Furthermore, in a fully unsupervised setting, we can separate accurate from inaccurate predictions with an F1 score of 0.666, demonstrating that consistency-based signals contain information for error detection without any labeled data.
[…]
fig:roc_comparison
Figure 4: Using signals from self- and cross-model consistency provides high AUC for accuracy classification. Proxy benchmark results do not provide significant prediction signal as compared to consistency. Model AUCs range from 0.703 to 0.981, indicating strong prediction capability. LLaMA-3.3-70B-Instruct’s high AUC can be attributed to its low performance on the given task, creating an arbitrary classification problem and highlighting limitations of including lower-performing models in analyses.
5 Discussion
As LLMs become increasingly capable across tasks, understanding their limitations and quantifying accuracy for specific applications is critical for deployment (Unell et al., 2025). This work makes three main contributions in the high-stakes medical domain. First, we formalize guideline-concordant treatment recommendation generation and introduce a dataset of 121 patient cases annotated by board-certified oncologists. Second, we compare synthetic supervision and consistency-based supervision for benchmarking models on tasks lacking human annotations, finding that consistency-based benchmarks best approximate human judgments (Spearman $r=0.88$, RMSE=0.08). Third, we develop a hybrid agent and meta-classifier trained using unsupervised features, achieving an average of 0.800 AUC on held-out models and producing confidence scores that support ROC analysis, error trade-off optimization, and regulatory compliance (U.S. Food and Drug Administration, 2022). Feature analysis shows self-consistency and cross-model consistency are the strongest predictors of recommendation accuracy.
Several promising directions emerge from this work. Expanding the dataset to other cancer types and guidelines would help characterize the generalizability of our consistency-based benchmarking approach, potentially enabling cross-domain accuracy prediction (Sorin et al., 2024; Pool and Trujillo, 2024). Further evaluation of self- and cross-model consistency as a signal for accuracy classification across broader model sizes and architectures could deepen understanding of this method’s robustness. Increasing both the size of the human-annotated dataset and the number of dually annotated samples would strengthen confidence in ground-truth labels. Since clinicians exhibit variability in path selection, future work should also explore ways to explicitly model and incorporate human uncertainty into evaluation. Adaptive learning approaches that account for uncertainty across clinical scenarios may further improve sensitivity of proxy metrics (Kompa et al., 2021; Jing et al., 2024). Finally, investigating how proxy benchmark data can be leveraged for alignment with downstream human preferences could help mitigate data bottlenecks that limit current methods (Kim and Seo, 2024; Lee et al., 2025).
Appendix A Annotation Details
We begin by extracting the decision tree pages from the NCCN NSCLC guidelines (National Comprehensive Cancer Network, 2025) and creating a JSON-based representation where each decision point is labeled as a ”node.” Terminal nodes with ”recommendation” labels contain final treatment recommendations for clinicians, while intermediate nodes represent decision points requiring additional clinical information.
To evaluate model capability in mapping these guidelines onto real clinical data, we created a human-annotated dataset for guideline-adherent reasoning in NSCLC. Expert annotators traced patient notes through the NCCN guideline decision tree, recording the ordered sequence of page–node identifiers (e.g., NSCL-1-1 → NSCL-2-1 → … → terminal). Each annotation aimed to reach either the appropriate ”recommendation” node when sufficient clinical information was available, or the furthest node that could be accurately determined given the available patient data. Annotators additionally marked whether the observed care was guideline-compliant and provided brief rationales for any identified as non-compliance.
Thirteen physicians performed the annotations: twelve board-certified oncologists and one hematology–oncology fellow, averaging 13 years of clinical experience. Participants were recruited through data vendors, screened for oncology knowledge and rubric literacy, and verified via official certification registries. All physicians were compensated, and annotation instructions are provided.
A subset of cases was double-annotated, and disagreements were resolved through adjudication meetings, with all decisions documented for consistency. A strict no-AI-use policy was enforced via Insightful time-tracking with periodic screenshot audits (Insightful, 2025). Annotation time averaged 46.5 minutes per case. All cases were fully de-identified and contained no protected health information, so the study did not constitute human-subjects research and no IRB was required.
Clinical Annotation Instructions You are helping us understand how a synthetic patient progressed through the NCCN guidelines for non-small cell lung cancer (NSCLC), based on information in their clinical note. Your task is to carefully read the clinical note and determine which steps in the NCCN guideline the patient appears to have gone through. Using your clinical judgment and the structure of the guideline, trace the path the patient has followed so far. Path Format: You will identify the treatment path in the following format: NSCL-x-y $\rightarrow$ NSCL-a-b $\rightarrow\rightarrow$ Final Treatment Node Each step should include the full node ID from the guideline (e.g., NSCL-1-1). Do not skip any nodes, even if some are broader or less specific than others. The final step should be a terminal node—a node that includes a clear treatment recommendation. If the patient’s treatment has not reached a terminal node, you should provide the recommended pathway to the next terminal node, as permitted by the information in the note. Compliance Assessment: There is a box to check if the patient’s path complies with the NCCN guidelines. If you believe the patient’s path does not align with the NCCN guideline, please do not check the box. In this case, you will also include a brief explanation of why the treatment path doesn’t match the guidelines. Only include a “reason” if the path is not compliant. If the path is guideline-compliant, just provide the full treatment path. Important Notes: You do not need to speculate beyond what is described in the note You do not need to assume the patient has already received the final treatment—just determine the most appropriate next step or current position in the guideline In some cases, the NSCLC Guidelines may not apply. For these cases, please note that the guidelines do not apply and proceed to the next task Materials Provided: A clinical note (the patient’s case) The full NCCN guideline (for reference) Tips and Tricks: We recommend utilizing “Command + f” (Mac) or “Ctrl+f” (Windows) to find key terms and parts of the note/guideline tree that are relevant. The guidelines consist of many “pages” (i.e. NSCLC-1, NSCLC-2) which contain further questions that will bring you to other pages. Example Workflow: Step 1: REDACTED DUE TO NCCN LICENSING REGULATIONS Step 2: Using Ctrl+F (Command+F on Mac), proceed to the next step based on the patient’s clinical notes. Per the example above, search “NSCL-2” to skip to the section beginning: ‘‘NSCL-2’’: { ‘‘page_id’’: ‘‘NSCL-2’’ Under NSCL-2, you see “nodes” and then a numbered list of different criteria. Begin with node 1 and follow the branching guidance from there. At each step, confirm that the treatment outlined in the patient’s clinical note is in accordance with the NCCN Guidelines. If the treatment diverges from the path outlined by the guidelines, it would be considered noncompliant. The treatment path you provide should be the treatment path that was followed in the patient’s care. If this differs from the path recommended by the NCCN Guidelines, indicate that the path was noncompliant.
Appendix B Path Overlap Score
Let $\mathcal{P}=\{P_{1},P_{2},\ldots,P_{k}\}$ be a collection of $k$ predicted paths, where
$P_{i}=(p_{i,1},p_{i,2},\ldots,p_{i,n_{i}}),\quad i=1,\ldots,k,$
and $p_{i,j}$ denotes the $j$-th node in path $P_{i}$.
Define the node set of path $P_{i}$ as
$V_{i}=\{p_{i,1},p_{i,2},\ldots,p_{i,n_{i}}\},$
and the union and intersection of all paths as
$U=\bigcup_{i=1}^{k}V_{i},\quad I=\bigcap_{i=1}^{k}V_{i}.$
The Path Overlap Score (corresponding to path_match_fraction in code) is defined as the Jaccard similarity:
$\text{Overlap}(\mathcal{P})=\begin{cases}\frac{|I|}{|U|},&|U|>0,\\ 1,&|U|=0.\end{cases}$
Properties
  • $\text{Overlap}(\mathcal{P})\in[0,1]$.
  • $\text{Overlap}(\mathcal{P})=1$ if and only if all paths share exactly the same set of nodes.
  • Higher scores indicate greater similarity in node coverage across paths.
  • Symmetric with respect to path ordering.
  • Sensitive only to node membership, not ordering or repetition.
Appendix C Treatment Match
Final Treatment Consistency Score
Let $\mathcal{P}=\{P_{1},P_{2},\ldots,P_{k}\}$ be a set of $k$ predicted treatment paths for a given patient, where each path $P_{i}=(p_{i,1},p_{i,2},\ldots,p_{i,n_{i}})$ terminates with a final treatment recommendation $f_{i}=p_{i,n_{i}}$.
Case 1: Ground Truth Available
When ground truth final treatment $f^{*}$ is available, the score is computed as the indicator function:
$\mathbf{1}_{f_{i}=f^{*}}=\begin{cases}1&\text{if }f_{i}=f^{*}\\ 0&\text{otherwise}\end{cases}$
This returns whether there was an exact match between the ground truth and the prediction final treatment.
Case 2: No Ground Truth Available
In the absence of ground truth, we measure internal consistency by computing the proportion of repeated final treatments:
$S_{\text{final}}(\mathcal{P})\;=\;\frac{\max_{t\in T}c(t)}{k}$
where:
$k$ = $=\text{ total rollouts}$ (1)
$c(t)$ $=\sum_{i=1}^{k}\mathbf{1}_{f_{i}=t}\quad\text{(count of paths with final treatment }t\text{)}$ (2)
The numerator $\max_{t\in T}c(t)$ counts the total number of ”repeated” final treatment occurrences (i.e., how many times each treatment appears beyond its first occurrence) and returns the frequency of the mode treatment selection. For all models $k=10$.
Properties
  • $S_{\text{final}}(\mathcal{P})\in[0,1]$ for both cases
  • Higher scores indicate better accuracy (Case 1) or higher consensus (Case 2)
  • Case 2: Score of 0 indicates complete disagreement; score approaching 1 indicates strong consensus
Appendix D Error Analysis
We find that 40.42% of all errors made by models on human-annotated clinical text are flagged within the top 5 points of confusion by the model itself during its internal rollouts. This enables practitioners to refine instructions, perform fine-tuning, or apply other interventions to address these frequent errors—without relying on human annotations to detect them. Discrepancies mostly arise around guideline compliance and tumor staging, indicating that models struggle to differentiate between non-compliant and compliant cases with high accuracy as well as are unable to leverage parametric knowledge to perform tumor staging. Potential interventions could be training a model specifically for either task or providing additional context with clear instructions regarding clinical expectations.
[…]
fig:ea_comparison
Figure 7: Identification of most common discrepancies between human annotations and model annotations compared to most common discrepancies between $k$ model rollouts of path prediction.
Appendix E Consistency-Accuracy Correlation
We show that for both Path Overlap metric and Treatment Match metrics, consistency of a prediction is correlated with accuracy. We do see, however, that there are different levels of correlation, indicating that consistency-based approaches inherently are more accurate for some model families than others. We note that DeepSeek-R1 has a low outlier correlation coefficient value. This highlights the robustness of the meta-classification pipeline to varying model internal consistencies. Even though there is no direct correlation between self-consistency and accuracy for this model, we are still able to classify accuracy with high fidelity in our method.
Table 3: Iteration consistency vs. accuracy correlation ($r$) for path overlap and treatment match. Mean $\pm$ SEM across models are reported at the bottom.
Model Path Overlap Treatment Match
GPT-5-High 0.477 0.472
GPT-4.1 0.812 0.923
GPT-5-Minimal 0.925 0.866
o3 0.700 0.795
o4-mini 0.566 0.491
GPT-5-Medium 0.794 0.935
DeepSeek-R1 -0.647 0.129
LLaMA-3.3-70B-Instr. 0.688 0.789
Mean $\pm$ SEM 0.540 $\pm$ 0.179 0.675 $\pm$ 0.103
Appendix F Unsupervised clustering performance
We cluster on features related to self- and cross-model consistency. We are able to distinguish between True Negatives and False Negatives with high accuracy, and are able to derive some signal regarding the True Positives. This indicates that consistency alone can be used as a baseline in an unsupervised method to identify accuracy of prediction. The performance is below supervised approaches, as to be expected, but is promising for further applications in which supervision is not viable or possible. F1 score is .666, indicating potential of unsupervised methods in conjunction with consistency approaches to evaluate model performance in zero-label settings.
Figure 16: Confusion Matrix for K-means clustering over consistency-derived features.
Appendix G Meta-Classifier Performance on Internal Features
We report the performance of our meta-classifier using internal features only for training. This enables us to disaggregate classification performance between cross-model and self consistency. We see here that the AUROC scores for each model is lower than when we include cross-model information, but it is still significantly above random, indicating that while cross-model consistency is a useful feature, it is not the only feature that allows for insight into model performance.
Figure 17: ROCs disaggregated by model for classification of treatment prediction accuracy using exclusively internal features during training.
Appendix H Model Information
Table 4: Model release dates and API versions.
Model Provider API Version
DeepSeek-R1 Azure AI 2024-05-01-preview
LLaMA-3.3-70B-Instr. Azure OpenAI 5 (Model release: 2024-12-06)
GPT-4.1 Azure OpenAI 2024-12-01-preview
GPT-5 Azure OpenAI 2024-12-01-preview
o4-mini Azure OpenAI 2024-12-01-preview
o3 Azure OpenAI 2024-12-01-preview
For supervised classification, we trained a logistic regression classifier with L2 regularization (C=0.1), optimized using LBFGS with up to 10,000 iterations, and applied class-balanced weighting to account for label imbalance.